OpenStudy (tisispurecuckery):

How can the properties of exponents help solve exponential equations?

OpenStudy (tisispurecuckery):

@jh99

OpenStudy (jh99):

exponent and logarithmic have the sanme base

OpenStudy (tisispurecuckery):

How can the properties of exponents help solve logarithmic equations?

OpenStudy (jh99):

I need a specific example for this one. \[b=\log(a) \] \[10^b=a\] are the same thing

OpenStudy (tisispurecuckery):

That's all the question says.

OpenStudy (jh99):

\[\log_{b}a=c \] is the same as \[b^c=a\]

OpenStudy (jh99):

Easy way to remember this: \[\log_{10}100=2 \]\[10^2=100\]